LA’s Sports Equipment began business in 2019, and its sales in 2019 were 4,800 units of sports equipment. The selling price per unit was $300. Variable cost per unit of sports equipment was $186. Annual fixed costs were $446,400.

1. Please use the format of a contribution margin income statement to calculate the contribution margin (in dollars), unit contribution, and contribution margin ratio for LA’s Sports Equipment. Total Per Unit Percentage Sales (xxx units), Less: Variable costs, Contribution margin, Less: Fixed costs Net income

2. What will be the break-even point in units and in dollars if we assume that LA’s Sports Equipment sells one type of sports equipment only?

3. How many units must the company sell to earn a pre-tax income of $500,028?

4. If the company’s tax rate is 28 percent, how many units of sports equipment must be sold to earn an after-tax profit of $700,200?

5. What is the company’s margin of safety based on the sales of 4,800 units of sports equipment?

6. Assume that there is no change in fixed costs. What will be the expected net income if this company has the sales revenue of $2,010,000?
Answer #1

Answer to Requirement 1:

Per Unit 300.00 186.00 Sales (4,800 units) Less: Variable Cost Contribution Margin Less: Fixed Cost Net Income Total $ 1,440,

Answer to Requirement 2:

Breakeven Point in units = Fixed Costs / Contribution Margin per unit
Breakeven Point in units = $446,400 / $114
Breakeven Point in units = 3,916

Breakeven Point in dollar = Fixed Costs / Contribution Margin Ratio
Breakeven Point in dollar = $446,400 / 0.38
Breakeven Point in dollar = $1,174,737

Answer to Requirement 3:

Required Sales in units = (Fixed Costs + Pre-tax Income) / Contribution Margin per unit
Required Sales in units = ($446,400 + $500,028) / $114
Required Sales in units = 8,302

Answer to Requirement 4:

After-tax Profit = Pre-tax Income * (1 - Tax Rate)
$700,200 = Pre-tax Income * (1 - 0.28)
Pre-tax Income = $972,500

Required Sales in units = (Fixed Costs + Pre-tax Income) / Contribution Margin per unit
Required Sales in units = ($446,400 + $972,500) / $114
Required Sales in units = 12,446

Answer to Requirement 5:

Actual Sales = Units Sold * Selling Price per unit
Actual Sales = 4,800 * $300
Actual Sales = $1,440,000

Margin of Safety = Actual Sales - Breakeven Point in dollar
Margin of Safety = $1,440,000 - $1,174,737
Margin of Safety = $265,263

Answer to Requirement 6:

Pre-tax Income = Sales * Contribution Margin Ratio - Fixed Costs
Pre-tax Income = $2,010,000 * 38.00% - $446,400
Pre-tax Income = $317,400

After-tax Profit = Pre-tax Income * (1 - Tax Rate)
After-tax Profit = $317,400 * (1 - 0.28)
After-tax Profit = $228,528